A client has tetrachloroethylene (PCE) contaminated groundwater at a site. It is possible to use PCE as an electron acceptor, resulting in TCE. The sequence of anaerobic reduction for chlorinated solvents is:

PCE reduction: ½ Cl2C=CCl2 + ½ H+ + e- → ½ Cl2C=CHCl + ½ Cl- ΔG0’ = -41.42 kJ/e- eq

TCE reduction: ½ Cl2C=CHCl + ½ H+ + e- → ½ ClHC=CHCl + ½ Cl- ΔG0’ = -40.17 kJ/e- eq

DCE reduction: ½ ClHC=CHCl + ½ H+ + e- → ½ ClHC=CH2 + ½ Cl- ΔG0’ = -30.12 kJ/e- eq

VC reduction: ½ ClHC=CH2 + ½ H+ + e- → ½ H2C=CH2 + ½ Cl- ΔG0’ = -36.82 kJ/e- eq

Acetate is proposed to be used as an electron donor to remediate the PCE (O-1 in Table 2.3). Assuming ε = 0.75, determine whether acetate will provide sufficient energy to cause the PCE to be fully degraded to ethylene. The acetate would be dissolved in water and pumped in underground with a fertilizer solution including ammonium and phosphate.
Answer #1

do ta Lauat According to the following data, - Addih all the equations. 1/2 ell=cela +2494 40 1/240 = Clutpelo. 0643198.53€3l